Based on the following information, calculate the sustainable growth rate for Kaleb’s Heavy Equipment:
  Profit margin 9.8%
  Capital intensity ratio .49    
  Debt-equity ratio .57    
  Net income $23,000    
  Dividends $15,870    

Multiple Choice

  • 8.44%

  • 7.88%

  • 27.76%

  • 10.78%

  • 11.28%
Answer #1

Answer: Correct answer is 10.78%

Profit margin=Net income/Sales
=>Sales=Net income/Profit margin=23000/9.8%=234693.8776

Capital intensity ratio=Total assets/Sales
=>Total assets=Capital intensity ratio*Sales=.49*234693.8776=115000

Debt/Equity=.57
=>Debt=.57*Equity
Total assets=Debt+Equity=.57*Equity+Equity
=>Total assets=1.57Equity
=>115000=1.57Equity
=>Equity=115000/1.57=73248.40764

Dividend payout ratio=Dividends/Net income=15870/23000=0.69

Retention ratio=1-Dividend payout ratio=1-0.69=0.31

Return on equity (ROE)=Net income/Equity=23000/73248.40764=0.314

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]
=(0.314*0.31)/[1-(0.314*0.31)]
=0.09734/0.90266
Sustainable growth rate=0.107836838 or 10.78% (Rounded to two decial places)